Bertha Co purchased equipment 4 years ago for $1,663,081. The equipment was depreciated straight line to zero over 7 years. What’s the current accumulated depreciation for the equipment?
Answer #1

Annual depreciation=(Cost-Salvage value)/Useful Life

=1,663,081/7=$237583/year

Hence current accumulated depreciation=(237583*4)

=$950332